### v1

@@ 6. Stability Analysis via ${\mathcal{K}}$-Lorentzian polynomials
Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k},d\geq 5$ be a univariate polynomial with positive coefficients and $a_{k}^{2}>x_{0}a_{k-1}a_{k+1}$ where $x_{0}$ is the unique positive root of the polynomial $x^{3}-x^{2}-2x-1(x_{0}\approx 2.1479)$. Then
  1. (a)
    $f(t)$ is Hurwitz-stable and
  2. (b)
    if the characteristic polynomial $\chi(A)=f(t)$, the corresponding dynamical system $\frac{dx}{dt}=Ax$ is stable.
@@
Let $f(x)$ be a strictly ${\mathcal{K}}$-CLC of degree $d\leq 4$ with $v\in\operatorname{int}{\mathcal{K}}$. If the characteristic polynomial $\chi(A)=f(x_{0}+tv)$ for some $x_{0}\in\operatorname{int}{\mathcal{K}}$, the dynamical system $\frac{dx}{dt}=Ax$ is stable.
@@
If $f(x)$ is a strictly ${\mathcal{K}}$-CLC of degree $d\leq 4$, then by Theorem 4.20, $f(x)$ is Hurwitz-stable polynomial over ${\mathcal{K}}$. Then it follows from Definition 4.5 that univariate restriction $f(x+tv)$ is Hurwitz-stable for any $x,v\in\operatorname{int}{\mathcal{K}}$. Thus, the matrix $A$ is (Lyapunov-type) stable if $\chi(A)=f(x_{0}+tv)$ for $x_{0},v\in\operatorname{int}{\mathcal{K}}$ by Theorem 4.2. Equivalently, all the leading principal minors of the corresponding Hurwitz matrix ${\mathrm{Hur}}_{f(x_{0}+tv)}$ are positive for any $x_{0},v\in\operatorname{int}{\mathcal{K}}$. ∎
@@
Let $f(x)$ be a strictly ${\mathcal{K}}$-CLC of degree $d\geq 5$ with $v\in\operatorname{int}{\mathcal{K}}$. If the characteristic polynomial $\chi(A)=f(x_{0}+tv)=\sum_{k=0}^{d}a_{k}t^{k}$ with $a_{k}^{2}>\alpha a_{k-1}a_{k+1}$ where $\alpha$ is the unique positive root of the polynomial $t^{3}-t^{2}-2t-1(\alpha\approx 2.1479)$ for any $x_{0}\in\operatorname{int}{\mathcal{K}}$, the dynamical system $\frac{dx}{dt}=Ax$ is stable.
6.1. Cone-Stability
Consider the linear time invariant system (IVP)
$\frac{dx(t)}{dt}+Ax(t)=0,x(0)=x_{0}\in{\mathcal{K}}.$
For example, in $2D$ system, consider $A=\begin{bmatrix}1&2\\ 1&1\end{bmatrix}$ and ${\mathcal{K}}={\mathbb{R}}^{2}_{\geq 0}$, cf. [GB04, Example 1], [STH22, Example 1]. Clearly, $A$ is not a stable matrix, since $\operatorname{eig}(A)=\{1+\sqrt{2},1-\sqrt{2}\}$. Another example in 3D system, let $A=\begin{bmatrix}1&3&2\\ 5&-1&1\\ -3&10&2\end{bmatrix}$ and ${\mathcal{K}}={\mathbb{R}}^{3}_{\geq 0}$. $A$ is not a stable matrix, since $\operatorname{eig}(A)=\{6.2773,-2.1387+1.5087i,-2.1387-1.5087i\}$. Thus, the origin is unstable when systems are not constrained on ${\mathcal{K}}$. It is shown in [GB04], [STH22] that the trivial solution, 00 is asymptotically stable on ${\mathcal{K}}={\mathbb{R}}^{2}_{\geq 0}$ and ${\mathbb{R}}^{3}_{\geq 0}$ respectively. In essence, the equilibrium point may not be stable in the entire state space, but is stable when the system states are constrained to a cone ${\mathcal{K}}$. Inspired by the evidence of the instances that constraints make the system stable even though the unconstrained system was not, we consider the following class of dynamical systems as outlined in [GB04], [STH22], and references therein.
The notation $C^{0}([t_{0},+\infty);{\mathbb{R}}^{d})$ represents the space of continuous functions defined on the interval $[t_{0},+\infty)$,and taking values in ${\mathbb{R}}^{d}$. The notation $L^{\infty}_{\mathrm{loc}}(t_{0},+\infty;{\mathbb{R}}^{d})$ represents the space of locally essentially bounded functions defined on $(t_{0},+\infty)$ with values in ${\mathbb{R}}^{d}$. The notation $C^{1}({\mathbb{R}}^{d};{\mathbb{R}})$ represents the space of continuously differentiable scalar-valued functions defined on ${\mathbb{R}}^{d}$.
The class of Dynamical Systems: Let ${\mathcal{K}}\subset{\mathbb{R}}^{d}$ be a nonempty closed convex set. Let $A\in{\mathbb{R}}^{d\times d}$ be a given matrix and $F:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ a nonlinear operator. For $(t_{0},x_{0})\in{\mathbb{R}}\times{\mathcal{K}}$, we consider the following problem $P(t_{0},x_{0})$: Find a function $t\rightarrow x(t)(t\geq t_{0})$ with $x\in C^{0}([t_{0},+\infty);{\mathbb{R}}^{d})$ , and $\frac{dx}{dt}\in L^{\infty}_{\mathrm{loc}}(t_{0},+\infty;{\mathbb{R}}^{d})$ such that
(2) $\begin{cases}&\langle\frac{dx}{dt}(t)+Ax(t)+F(x(t)),v-x(t)\rangle\geq 0,\ \ \forall v\in{\mathcal{K}},\ \mbox{a.e}\ t\geq t_{0}\\ &x(t)\in{\mathcal{K}},\ \ \ \ t\geq t_{0}\end{cases}$
$x(t_{0})=x_{0}.$
This is known as an evolution variational inequality (EVI) dynamical system. Recall that the normal cone of a closed convex set ${\mathcal{K}}$ at a point $x\in{\mathcal{K}}$ is defined as:
$N_{{\mathcal{K}}}(x)=\{y\in{\mathbb{R}}^{n}:\langle y,v-x\rangle\leq 0,\forall v\in{\mathcal{K}}\}$
The tangent cone is the polar of the normal cone:
$T_{{\mathcal{K}}}(x)=\{d\in{\mathbb{R}}^{n}:\langle d,y\rangle\leq 0,\forall y\in N_{{\mathcal{K}}}\}=\mathrm{cl}\{d\in{\mathbb{R}}^{n}:d=\alpha(v-x),v\in{\mathcal{K}},\alpha\geq 0\}$
Then, using the standard convex analysis tools one can equivalently rewrite the EVI dynamical system $P(t_{0},x_{0})$ as
$\begin{cases}\langle\frac{dx}{dt}+Ax+F(x),v-x\rangle\geq 0,\ \forall v\in{\mathcal{K}},\ \mbox{a.e}\ t\geq t_{0}\\ x(t)\in{\mathcal{K}},\ \ t\geq t_{0}\end{cases}\Leftrightarrow\begin{cases}\frac{dx}{dt}+Ax+F(x)\in-N_{{\mathcal{K}}}(x)\\ x(t)\in{\mathcal{K}},t\geq t_{0}\end{cases}$
Here we mention the conditions for the existence and uniqueness of the initial value problem $P(t_{0},x_{0})$ which is a special case of more general setting outlined in [GMM03, Theorem 2.1-Kato, Cor 2.2].
Theorem 6.7.
[GB04] Let ${\mathcal{K}}\subset{\mathbb{R}}^{d}$ be a nonempty closed convex set of ${\mathbb{R}}^{d}$ and let $A\in{\mathbb{R}}^{d\times d}$ be a real matrix of order $d$. Suppose that $F:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ can be written as
$F=F_{1}+\Phi^{{}^{\prime}}$
where $F_{1}$ is Lipschitz continuous and $\Phi\in C^{1}({\mathbb{R}}^{d};{\mathbb{R}})$ is convex. Let $t_{0}\in{\mathbb{R}}$ and $x_{0}\in{\mathcal{K}}$ be given. Then there exists a unique $x(t):=x(.;t_{0},x_{0})\in C^{0}([t_{0},+\infty);{\mathbb{R}}^{d})$, such that $\frac{dx}{dt}\in L^{\infty}_{\mathrm{loc}}(t_{0},+\infty;{\mathbb{R}}^{d})$, $x$ is right differentiable on $[t_{0},+\infty)$, and the conditions of (2) are satisfied.
@@ 6.1. Cone-Stability
Additionally, if it is assumed that $0\in{\mathcal{K}}$ and $\langle F(0),v\rangle\geq 0$ for all $v\in{\mathcal{K}}$, then the unique solution $x(t,t_{0},0)=0$, $t\geq t_{0}$, i.e., the trivial solution 00 is the unique solution of the problem $P(t_{0},0)$. We consider the class of dynamical system which satisfies all the conditions mentioned in Theorem 6.7 along with $0\in{\mathcal{K}}$ and $\langle F(0),v\rangle\geq 0$ for all $v\in{\mathcal{K}}$. We refer such systems as $P(t_{0},x_{0})$ throughout the rest of the paper unless stated otherwise.
We now define stability of an equilibrium point w.r.t a closed convex set ${\mathcal{K}}$ for the EVI system as outlined in [GB04], [STH22].
Definition 6.9.
The equilibrium point $x=0\in{\mathcal{K}}$ is called (Lyapunov-type) stable w.r.t ${\mathcal{K}}$ if for any $\epsilon>0$, there exists a $\delta>0$ such that for any initial condition $x_{0}\in{\mathcal{K}}$ with $\|x_{0}\|\leq\delta$, the solution $x(t)\in{\mathcal{K}}$ of the system EVI satisfies $\|x(t)\|<\epsilon$ for all $t\geq t_{0}$. The equilibrium point $x=0\in{\mathcal{K}}$ is called asymptotically stable w.r.t ${\mathcal{K}}$ if it is stable and there exists a $\delta>0$ such that for any $x_{0}\in{\mathcal{K}}$ with $\|x_{0}\|\leq\delta$, the solution $x(t)\in{\mathcal{K}}$ of the EVI system approaches to the origin, i.e., $\lim_{t\rightarrow\infty}\|x(t)\|=0$, equivalently, $\lim_{t\rightarrow\infty}x(t)=0$.
This means that any trajectory within the cone that starts sufficiently close to 00 remains within a specified distance from 00 for all future times. In this paper, we assume that ${\mathcal{K}}$ is a proper convex cone. However, these results also apply to any closed convex set ${\mathcal{K}}$ that contains the origin; see [GB04] for details. Here we discuss some results related to ${\mathcal{K}}$-copositive matrices, stability w.r.t ${\mathcal{K}}$, and strictly ${\mathcal{K}}$-CLC. In order to do that we first report some general abstract theorems of stability, asymptotic stability and instability w.r.t ${\mathcal{K}}$ in terms of generalized Lyapunov functions, cf. [GB04].
Theorem 6.10.
Consider a system $P(t_{0},x_{0})$. Suppose there exist $\sigma>0$ and $V\in C^{1}({\mathbb{R}}^{d};{\mathbb{R}})$ such that
  1. (1)
    $V(x)\geq a(\|x\|),\ x\in{\mathcal{K}},\|x\|\leq\sigma$ with $a:[0,\sigma]\rightarrow{\mathbb{R}}$ satisfying $a(t)>0$, for all $t\in(0,\sigma)$;
  2. (2)
    $V(0)=0$;
  3. (3)
    $x-V^{{}^{\prime}}(x)\in{\mathcal{K}}$, for any $x\in\partial{\mathcal{K}}$ and $\|x\|\leq\sigma$ ;
  4. (4)
    $\langle Ax+F(x),V^{{}^{\prime}}(x)\rangle\geq 0$, for all $x\in{\mathcal{K}}$ and $\|x\|\leq\sigma$.
Then, the trivial solution of $P(t_{0},x_{0})$ is stable w.r.t ${\mathcal{K}}$.
Theorem 6.11.
Consider a system $P(t_{0},x_{0})$. Suppose there exist $\lambda>0,\sigma>0$ and $V\in C^{1}({\mathbb{R}}^{d};{\mathbb{R}})$ such that
  1. (1)
    $V(x)\geq a(\|x\|),\ x\in{\mathcal{K}},\|x\|\leq\sigma$ with $a:[0,\sigma]\rightarrow{\mathbb{R}}$ satisfying $a(t)>ct^{\tau}$, for all $t\in[0,\sigma]$, for some constants $c>0,\tau>0$;
  2. (2)
    $V(0)=0$;
  3. (3)
    $x-V^{{}^{\prime}}(x)\in{\mathcal{K}}$, for any $x\in\partial{\mathcal{K}}$ and $\|x\|\leq\sigma$ ;
  4. (4)
    $\langle Ax+F(x),V^{{}^{\prime}}(x)\rangle\geq\lambda V(x)$, for all $x\in{\mathcal{K}}$ and $\|x\|\leq\sigma$.
Then, the trivial solution of $P(t_{0},x_{0})$ is asympotically stable w.r.t ${\mathcal{K}}$.
Remark 6.12.
The condition $(3)$ is a sufficient condition which implies $-V^{{}^{\prime}}(x)\in T_{{\mathcal{K}}}(x)$ for all $x\in\partial{\mathcal{K}}$, $\|x\|\leq\sigma$, where $T_{{\mathcal{K}}}(x)$ is the tangent cone to ${\mathcal{K}}$.
We discuss Lyapunov Stability on ${\mathcal{K}}$ as outlined in [GB04].
Definition 6.13.
The matrix $A\in{\mathbb{R}}^{n\times n}$ is Lyapunov positive stable (semi-stable) on ${\mathcal{K}}$ if there exists a matrix $P\in{\mathbb{R}}^{n\times n}$ such that
  1. (a)
    $\mathrm{inf}_{x\in{\mathcal{K}}\setminus\{0\}}\frac{x^{t}Px}{\|x\|^{2}}>0$;
  2. (b)
    $\langle Ax,[P+P^{t}]x\rangle>0(\geq 0)$ for all $x\in{\mathcal{K}}$;
  3. (c)
    $x\in\partial{\mathcal{K}}\Rightarrow(I-[P+P^{t}])x\in{\mathcal{K}}$.
Let $\mathcal{P}_{{\mathcal{K}}}$ ($\mathcal{P}^{+}_{{\mathcal{K}}}$) denote the set of ${\mathcal{K}}$- copositive (resp. strictly copositive) matrices and
$\mathcal{P}^{++}_{{\mathcal{K}}}=\{P\in{\mathbb{R}}^{n\times n}:\mathrm{inf}_{x\in{\mathcal{K}}\setminus\{0\}}\frac{x^{t}Px}{\|x\|^{2}}>0\}.$
Let ${\mathcal{L}}_{k}$ denote the set of all Lyapunov semi-stable matrices on ${\mathcal{K}}$, i.e.,
${\mathcal{L}}_{{\mathcal{K}}}=\{A\in{\mathbb{R}}^{n\times n}:\exists P\in{\mathcal{P}}^{++}_{{\mathcal{K}}}\ \mbox{such that}\ (I-[P+P^{t}])\partial{\mathcal{K}}\subset{\mathcal{K}}\ \mbox{and}\ A^{t}P+PA\in{\mathcal{P}}_{{\mathcal{K}}}\}$
and ${\mathcal{L}}^{++}_{k}$ denote the set of all Lyapunov positive stable matrices on ${\mathcal{K}}$, i.e.,
${\mathcal{L}}^{++}_{{\mathcal{K}}}=\{A\in{\mathbb{R}}^{n\times n}:\exists P\in{\mathcal{P}}^{++}_{{\mathcal{K}}}\ \mbox{such that}\ (I-[P+P^{t}])\partial{\mathcal{K}}\subset{\mathcal{K}}\ \mbox{and}\ A^{t}P+PA\in{\mathcal{P}}^{++}_{{\mathcal{K}}}\}$
Remark 6.14.
Note that $\langle Ax,[P+P^{t}]x\rangle>0\Leftrightarrow A^{t}P+PA\in{\mathcal{P}}^{++}_{{\mathcal{K}}}$. Let $A\in{\mathbb{R}}^{n\times n}$ be a positive strictly-stable matrix, i.e., the real part of all its eigenvalues are positive. Then there exists a matrix $P$ such that the conditions $(a)$ and $(b)$ in Definition 6.13 are satisfied since by Lyapunov Theorem there exists a positive definite matrix $P\in{\mathbb{R}}^{n\times n}$ such that $A^{t}P+PA=Q$, a positive definite matrix and $\langle Ax,[P+P^{t}]x\rangle=\langle Qx,x\rangle>0$ for all $x\in{\mathbb{R}}^{d}\setminus\{0\}$.
Lemma 6.15.
[GMM03], [GB04] If ${\mathcal{K}}$ is a proper convex cone, ${\mathcal{P}}^{++}_{{\mathcal{K}}}={\mathcal{P}}^{+}_{{\mathcal{K}}}$.
Furthermore, [GB04, Theorem 5] establishes that in the special case of problem $P(t_{0},x_{0})$ with $F\equiv 0$ referred to as the linear evolution variational inequality (LEVI), Lyapunov stability of matrix $A$ with respect to ${\mathcal{K}}$ can determine the stability of the trivial solution of the $P(t_{0},x_{0})$, LEVI system. Consider LEVI system defined as follows.
(3) $\begin{cases}\langle\frac{dx}{dt}+Ax,v-x\rangle\geq 0,\ \forall v\in{\mathcal{K}},\ \mbox{a.e}\ t\geq t_{0}\\ x(t)\in{\mathcal{K}},\ \ t\geq t_{0}\end{cases}$
$x(t_{0})=x_{0}$
Theorem 6.16.
[GB04]
  1. (a)
    If $A$ is ${\mathcal{K}}$-copositive, $A$ is Lyapunov semi-stable
  2. (b)
    If $A$ is strictly ${\mathcal{K}}$-copositive, $A$ is Lyapunov positive stable
Proof.
(a) Let $P=\frac{1}{2}I$. Conditions $(a)$ and $(b)$ are straightforward. Additionally, $(A-[P+P^{t}])x=0\in{\mathcal{K}}$. Thus, $A\in L_{{\mathcal{K}}}$.
(b) The choice of $P=\frac{1}{2}I$ still works. ∎
Theorem 6.17.
[GB04, Theorem 5] Let ${\mathcal{K}}$ be a proper convex cone and consider the LEVI system defined in (3). Then
  1. (a)
    If $A\in{\mathcal{L}}_{{\mathcal{K}}}$, then the trivial solution of the LEVI system is stable w.r.t ${\mathcal{K}}$.
  2. (b)
    If $A\in{\mathcal{L}}_{{\mathcal{K}}}^{++}$, then the trivial solution of the LEVI system is asymptotically stable w.r.t. ${\mathcal{K}}$.
Theorem 6.18.
If the quadratic form $x^{t}Ax$ is (strictly)-${\mathcal{K}}$-CLC (equivalently, (strictly) ${\mathcal{K}}$-Lorentzian), then the trivial solution of the LEVI system defined in (3) is (asymptotically) stable w.r.t ${\mathcal{K}}$.
Proof.
If the quadratic form $x^{t}Ax$ is ${\mathcal{K}}$-CLC or strictly ${\mathcal{K}}$-CLC, then the matrix $A$ is ${\mathcal{K}}$-copositive or strictly ${\mathcal{K}}$-copositive, respectively, as stated in Corollary 2.5. Additionally, if $A$ is ${\mathcal{K}}$-copositive or strictly ${\mathcal{K}}$-copositive, then $A$ is Lyapunov semi-stable or Lyapunov positive stable by Theorem 6.16. The claim then follows from Theorem 6.17. ∎
Corollary 6.19.
If the quadratic form $x^{t}Ax$ is (strictly)-${\mathcal{K}}$-CLC, then the trivial solution of the linear autonomous dynamical system of the form $\frac{dx}{dt}+Ax=0$ with $x(t)\in{\mathcal{K}}$ is (asymptotically) stable w.r.t ${\mathcal{K}}$.
Proof.
Note that the linear autonomous dynamical system $\frac{dx}{dt}+Ax=0$ is a special case contained in the general case, LEVI dynamical system. ∎
Here, we illustrate the situation through an example by visualizing the cone and system trajectories using MATLAB simulation.
Example 6.20.
Consider the quadratic form $q=-x_{1}^{2}+2x_{1}x_{2}-x_{2}^{2}+2x_{1}x_{3}+2x_{2}x_{3}-x_{3}^{2}$. It is not Lorentzian, but ${\mathcal{K}}$-Lorentzian where ${\mathcal{K}}={\mathcal{K}}(f,v)$ as shown in Figure 4. The corresponding system is $\frac{dx}{dt}+Ax=0$ where $A=\begin{bmatrix}-1&1&1\\ 1&-1&1\\ 1&1&-1\end{bmatrix}$. The MATLAB simulation code demonstrates how the trajectories converge to the origin when the initial conditions are selected from the cone.
Figure 3. ${\mathcal{K}}={\mathcal{K}}(f,v)$
Figure 4. Stability of the origin w.r.t ${\mathcal{K}}$
Corollary 6.21.
If the matrix $A$ is a (positive) nonnegative matrix with exactly one positive eigenvalue, the corresponding LEVI system defined in (3) is (asymptotically) stable over the nonnegative orthant.
Proof.
Note that $A$ is not a positive semidefinite matrix, so the system is not globally asymptotically stable. However, the corresponding quadratic form $x^{t}Ax$ is a (strictly) Lorentzian polynomial by Remark 2.8. Consequently, it follows from Theorem 6.18. ∎
Testing whether a quadratic form $q=x^{t}Qx$ is ${\mathcal{K}}$-CLC can be approached in several ways. The characterization of ${\mathcal{K}}$-CLC quadratic forms on three prominent self-dual cones is provided in [BD24]. For the nonnegative orthant (${\mathbb{R}}^{n}_{\geq 0})$, and the Lorentz cone ($\mathcal{L}_{n})$, verifying the ${\mathcal{K}}$-CLC property is relatively simpler compared to testing it over the positive semidefinite (PSD) cone. Let $q$ be a quadratic form on $\mathcal{S}^{n}$ whose matrix $Q$ is nonsingular and has exactly one positive eigenvalue. As shown in [BD24], $q$ is $\mathcal{S}^{n}_{+}$-Lorentzian if and only if the associated quartic polynomial $q(xx^{t})$ is nonnegative for all $x\in\mathbb{R}^{n}$. As a result, determining whether a quadratic form is $\mathcal{S}^{n}_{+}$-Lorentzian is a NP-hard problem, see [BD24] for details.
Now we discuss the case of nonlinear perturbations of LEVI systems. Let ${\mathcal{K}}$ be a proper convex cone. Consider the problem $P(t_{0},x_{0})$: Find $x\in C^{0}([t_{0},+\infty);{\mathbb{R}}^{d})$ where $\frac{dx}{dt}\in L^{\infty}_{\mathrm{loc}}(t_{0},+\infty;{\mathbb{R}}^{d})$ such that
(4) $\begin{cases}&x(t)\in{\mathcal{K}},\ \ \ \ t\geq t_{0}\\ &\langle\frac{dx}{dt}(t)+Ax(t)+F(x(t)),v-x(t)\rangle\geq 0,\ \ \forall v\in{\mathcal{K}},\ \mbox{a.e}\ t\geq t_{0}\\ \end{cases}$
$x(t_{0})=x_{0}$
Theorem 6.22.
[GB04, Theorem 7] Let ${\mathcal{K}}$ be a proper convex cone. Consider the dynamical system defined in (4), such that $-\Phi^{{}^{\prime}}(0)\in N_{{\mathcal{K}}}(0)$ and $\lim_{\|x\|\rightarrow 0}\frac{\|F_{1}(x)\|}{\|x\|}=0$ where $F=F_{1}+\Phi^{{}^{\prime}}$. If $A\in\mathcal{P}^{++}_{{\mathcal{K}}}$, then the trivial solution of this system is asymptotically stable.
Theorem 6.23.
If the quadratic form $x^{t}Ax$ is (strictly)-${\mathcal{K}}$-CLC, then the trivial solution of the nonlinear perturbed LEVI system defined in (4) is (asymptotically) stable w.r.t ${\mathcal{K}}$.
Proof.
Note that the assumptions $-\Phi^{{}^{\prime}}(0)\in N_{{\mathcal{K}}}(0)$ and $\lim_{\|x\|\rightarrow 0}\frac{\|F_{1}(x)\|}{\|x\|}=0$ ensure that $-F(0)\in N_{{\mathcal{K}}}(0)$, i.e., $\langle F(0),v\rangle\geq 0$ for all $v\in{\mathcal{K}}$. Therefore, the trivial solution 00 is the unique solution of $P(t_{0},0)$. If the quadratic form $x^{t}Ax$ is (strictly) ${\mathcal{K}}$-CLC, $A$ is (strictly) ${\mathcal{K}}$-copositive by Corollary 2.5. Then the result follows from Theorem 6.22
Acknowledgements. I am deeply grateful to Greg Blekherman and Marcella M. Gomez for their valuable suggestions on the earlier version of this article, which significantly contributed to its improvement.

### v2

@@ 5. Applications: Stability Analysis for LTI Systems
Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k},d\geq 5$ be a univariate polynomial with positive coefficients and $a_{k}^{2}>\sqrt{\alpha}a_{k-1}a_{k+1}$ where $\alpha$ is the unique positive root of the polynomial $t^{3}-t^{2}-2t-1(\alpha\approx 2.1479)$. Then the corresponding dynamical system $\frac{dx}{dt}=Ax$ is stable if the characteristic polynomial $\chi(A)=f(t)$,.
Proof.
It follows directly from Theorem 4.18 that $f(t)$ is Hurwitz-stable since the assumption $a_{k}^{2}>\sqrt{\alpha}\,a_{k-1}a_{k+1}$ for all $1\leq k\leq d-1$ implies the sufficient condition in that theorem. Then it follows from Theorem 5.1. ∎
@@
Let $f(x)$ be a strictly $\mathcal{K}$}-Lorentzian polynomial of degree $d\leq 4$ over a proper convex cone $\mathcal{K}$}. Suppose there exist $x_{0},v\in\operatorname{int}{\mathcal{K}}$ such that the characteristic polynomial of a matrix $A$ is given by
$\chi_{A}(t)=f(x_{0}+tv).$
Then the LTI system $\dot{x}=Ax$ is stable.
@@
By Theorem 4.19, $f$ is Hurwitz-stable over $\mathcal{K}$}, so for any $x,v\in\operatorname{int}{\mathcal{K}}$ the univariate restriction $t\mapsto f(x+tv)$ is Hurwitz-stable. In particular, $\chi_{A}(t)=f(x_{0}+tv)$ is Hurwitz-stable. The conclusion then follows from Theorem 5.1. ∎
@@
Let $f(x)$ be a strictly $\mathcal{K}$}-Lorentzian polynomial of degree $d\geq 5$ over a proper convex cone $\mathcal{K}$}. Suppose there exist $x_{0},v\in\operatorname{int}{\mathcal{K}}$ such that
$\chi_{A}(t)=f(x_{0}+tv)=\sum_{k=0}^{d}a_{k}t^{k}$
is the characteristic polynomial of a matrix $A$, and the coefficients satisfy
$a_{k}^{2}>\sqrt{\alpha}\,a_{k-1}a_{k+1},\qquad 1\leq k\leq d-1,$
where $\alpha$ is the unique positive root of $t^{3}-t^{2}-2t-1=0$ ($\alpha\approx 2.1479$). Then the LTI system $\dot{x}=Ax$ is stable.
Proof.
By the coefficient inequalities and Theorem 4.18, $\chi_{A}(t)$ is Hurwitz-stable. Stability of $\dot{x}=Ax$ then follows from Theorem 5.1. ∎
@@
Whenever the characteristic polynomial $\chi_{A}(t)$ of a stable matrix $A$ can be realized as a univariate restriction $\chi_{A}(t)=f(x_{0}+tv)$ of a Hurwitz-stable polynomial $f$ over a proper convex cone $\mathcal{K}$} with $x_{0},v\in\operatorname{int}{\mathcal{K}}$, we may regard $f$ as a $\mathcal{K}$}-Lorentzian lift of the dynamical system $\dot{x}=Ax$. This viewpoint connects LTI stability with the cone-geometric properties developed in this paper.
In particular, for a linear time-invariant system $\dot{x}=Ax$, Hurwitz stability of the characteristic polynomial $\chi_{A}(t)$ is equivalent to asymptotic stability of the equilibrium $x=0$. Thus, our Newton-type coefficient inequalities for strictly CLC and strictly $\mathcal{K}$}-Lorentzian polynomials yield simple sufficient tests ensuring stability of LTI systems directly from the coefficients of $\chi_{A}(t)$. Conversely, given a strictly $\mathcal{K}$}-CLC polynomial with suitable Hurwitz-stability properties, we can construct LTI systems whose spectral behaviour is encoded by $f$.
Acknowledgment: During the preparation of this manuscript, the author used ChatGPT to improve the clarity and conciseness of the writing. All content was subsequently reviewed, revised, and verified by the author, who takes full responsibility for the final version of the article.